OpenStudy (anonymous):

arithmetic a1=3 r=2 first 5 terms

OpenStudy (dumbcow):

An = A1+(n-1)*r

OpenStudy (anonymous):

explain please(:

OpenStudy (dumbcow):

thats the general formula, with that you can find any term A1 = 3 A2 = 3 + (2-1)*2 = 5 A3 = 3 +(3-1)*2 = 7 .... get the idea

OpenStudy (anonymous):

yah(: thanks!